OpenStudy (anonymous):

How do i find out the vertex of a parabola with the following information given: y=5x^2

OpenStudy (anonymous):

?

OpenStudy (mertsj):

Put the equation into vertex form which is y =a(x-k)^2 + k y = 5(x-0)^2+0 So we see that h is 0 and k is 0 and that is the vertex (h,k) or in your case (0,0)

OpenStudy (mertsj):

sorry I made a typo. y = a(x-h)^2 + k is the vertex form with vertex = (h,k)

OpenStudy (anonymous):

o

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

the vertex is (h,k) ???

OpenStudy (anonymous):

so h= x of vertex and k = y of vertex ????

OpenStudy (mertsj):

Well, you are supposed to put YOUR equation into that form. I will match them up for you again: y = a(x-h)^2+k y = 5(x-0)^2+0

OpenStudy (anonymous):

That's the vertex form of the quadratic equation

OpenStudy (anonymous):

so once i know \[h\] and\[k\] I USE THE \[h\] as an X VALUE FOR THE VERTEX and i use the \[k\] as a Y VALUE FOR THE VERTEX

OpenStudy (anonymous):

right?

OpenStudy (mertsj):

So a = 5, h = 0 and k = 0. so the vertex of the model equation is (h,k) but the vertex of your equation is the numbers that represent h and k in the specific equation which is your problem. Those numbers are h= 0 and k=0 so the vertex of y = 5x^2 is (0,0)

OpenStudy (mertsj):

BINGO!!!!! FAHMINATOR

OpenStudy (anonymous):

so every time the book asks me to find the vertex the easiest way to do it is to make it in y=a(x-h)^2+k FORM ?

OpenStudy (anonymous):

and h= x and k=y

OpenStudy (anonymous):

thank you so much my good man

OpenStudy (anonymous):

.

OpenStudy (mertsj):

I am not going to say that is the easiest way for every equation. In some cases it might be easier to use the fact that if the equation is in the form y = ax^2+bx+c, then it might be easier to say that the x value of the vertex is -b/2a. That would yield the same number as h. To find k you could plug that x value back into the equation and find the y value which would be the same as k. And I am not a man.

OpenStudy (anonymous):

The easiest way is to find the derivative of the quadratic equation. Then set it = to zero.

OpenStudy (mertsj):

Well that is not true if one does not know calculus.

OpenStudy (anonymous):

yeah i’m only in grade 10 i haven’t done that

OpenStudy (anonymous):

Same here grade 10